Perform the indicated operation and simplify.

(square root of 7 -5)(square root of 7 +5)

-8 + square root of 48 OVER 4

Answer by checkley71(8403)

SORRY I SKIPPED OVER THE SECOND PROBLEM.
(SQRT7-5)(SQRT7+5)=7-25=-18 ANSWER.
NOT SURE WHERE YOU NEED ().
-8+SQRT48/4=-8+SQRT12=-8SQRT(4*3)=-8*2SQRT3=-16SQRT2 ANSWER.
OR:
(-8+SQRT48)/4=[-8+SQRT(16*3)]4=(-8+4SQRT3)/4=-2+SQRT3 ANSWER.